# Amortized Inference of Multi-Modal Posteriors using Likelihood-Weighted Normalizing Flows

### v1

@@ 1 Introduction
Across diverse domains—from complex systems and finance to high-energy physics and astrophysics—scientific inquiry often relies on deriving theoretical parameters from observational data tarantola2005. At the core of this challenge lies the inverse problem: inferring the posterior distribution of theoretical parameters given a set of observables cranmer2020.
@@
Traditional approaches for posterior estimation rely on sampling algorithms such as Markov Chain Monte Carlo (MCMC) metropolis1953; hastings1970 and Nested Sampling (NS) skilling2006. In astrophysics and cosmology, implementations like emcee foreman2013 and dynesty speagle2020 have become standard tools. While these frameworks are statistically robust, they suffer significantly from the curse of dimensionality. In real-world scenarios, where the parameter space is high-dimensional and the likelihood function relies on computationally expensive simulators (e.g., in particle physics phenomenology brehmer2018), convergence can take weeks or even months.
@@
Recent advances in machine learning have introduced Normalizing Flows (NFs) as a powerful alternative for probabilistic modelling tabak2010; rezende2015. By learning a bijective mapping between a simple base distribution (e.g., a Gaussian) and the complex target distribution, NFs allow for exact density estimation and efficient sampling papamakarios2021 from the target distribution. Modern architectures, such as RealNVP dinh2016 and Neural Spline Flows durkan2019, offer enough expressivity to model highly complex distributions.
@@
However, standard training methodologies for NFs (typically Maximum Likelihood Estimation) rely on a critical assumption: that we possess a dataset of samples drawn from the target posterior papamakarios2017. In the context of Simulation-Based Inference (SBI) cranmer2020; lueckmann2021, this often leads to approaches like Neural Posterior Estimation (NPE) papamakarios2016; greenberg2019, which still require substantial simulation budgets to generate training data.
@@
In many scientific inverse problems, this assumption does not hold. We generally possess only a generic understanding of the parameters, hence assume they belong to a uniform distribution and a “black-box” simulator that can evaluate the likelihood of a given point. If we simply train a Normalizing Flow on samples drawn from the prior, the network will trivially learn to reconstruct the prior, failing to capture the information contained in the likelihood. While Variational Inference (VI) techniques exist to approximate posteriors blei2017, applying them efficiently with flows to capture multi-modal structures remains an active area of research wirnsberger2020; midgley2020.
@@
In this work, we propose a method to bridge this gap. By leveraging likelihood evaluations as sample weights during training—akin to Neural Importance Sampling muller2019—we demonstrate that a Normalizing Flow can directly learn the transformation from a simple base density to the complex target posterior, without ever seeing a “true” posterior sample. This results in an amortized inference framework that is computationally efficient and, crucially, capable of capturing the correct topological structure (modes) of the posterior distribution.
@@ 2.1 Normalizing Flows
Let $f_{\phi}:\mathcal{Z}\to\Theta$ be a differentiable bijection with a differentiable inverse (i.e., a diffeomorphism), parameterized by neural network weights $\phi$. We consider a latent space $\mathcal{Z}\subseteq\mathbb{R}^{D}$ endowed with a known, tractable base density $p_{Z}(z)$ (e.g., a standard Normal distribution or a Mixture of Gaussians), and a target parameter space $\Theta\subseteq\mathbb{R}^{D}$ associated with the unknown target distribution $p(\theta)$ (or the posterior $p(\theta|x)$).
The normalizing flow $f_{\phi}$ transforms the probability mass from the base distribution to the target. Crucially, because $f_{\phi}$ is a diffeomorphism, it preserves the topological properties of the domain. Consequently, the support and connectivity of the modelled distribution $q_{\phi}(\theta)$ are strictly constrained by the topology of the base density $p_{Z}(z)$.
@@
$q_{\phi}(\theta)=p_{Z}\left(z\right)\left|\det\!\left(\frac{\partial z}{\partial\theta}\right)\right|.$ (1)
@@
$q_{\phi}(\theta)$ $=p_{Z}\big(f_{\phi}^{-1}(\theta)\big)\,\left|\det\!\left(\frac{\partial f_{\phi}^{-1}(\theta)}{\partial\theta}\right)\right|.$ (2)
@@
$\log q_{\phi}(\theta)$ $=\log p_{Z}\big(f_{\phi}^{-1}(\theta)\big)+\log\left|\det\!\left(\frac{\partial f_{\phi}^{-1}(\theta)}{\partial\theta}\right)\right|.$ (3)
@@
2.2 Likelihood-Weighted Loss Function
The KL divergence from a reference probability distribution $r(\theta)$ to a model distribution $q_{\phi}(\theta)$ is defined as:
$KL\left(r(\theta)||q_{\phi}(\theta)\right)$ $=\int r(\theta)\log\frac{r(\theta)}{q_{\phi}(\theta)}d\theta$
@@ 2.2 Likelihood-Weighted Loss Function
$\mathcal{L}(\phi)=KL\left(r(\theta)\|q_{\phi}(\theta)\right)-\mathbb{E}_{\theta\sim r(\theta)}\left[\log(r(\theta))\right].$ (6)
Since the entropy term $\mathbb{E}\left[\log(r(\theta))\right]$ is independent of $\phi$, minimizing the negative log-likelihood is mathematically equivalent to minimizing the forward KL divergence:
$\min_{\phi}\left(\mathcal{L}(\phi)\right)\iff\min_{\phi}\left(KL\left(r(\theta)\|q_{\phi}(\theta)\right)\right).$ (7)
@@ 2.3 Application to Posterior Estimation
$KL\left(p(\theta\mid D)\|q_{\phi}(\theta)\right)$ $=\int p(\theta\mid D)\log\frac{p(\theta\mid D)}{q_{\phi}(\theta)}d\theta$
@@
$\min_{\phi}KL\left(p(\theta\mid D)\|q_{\phi}(\theta)\right)\iff\min_{\phi}\left(-\mathbb{E}_{\theta\sim p(\theta)}\left[\frac{p(D\mid\theta)}{Z}\log\left(q_{\phi}(\theta)\right)\right]\right).$ (10)
@@ 3 Implementation and Results
To evaluate the efficacy of the proposed likelihood-weighted training, we apply the method to a set of synthetic toy problems in 2 and 3 dimensions. These problems are designed with varying degrees of topological complexity (in terms of the number of modes) to test the expressivity of the flow. To quantify the fidelity of the posterior reconstruction, we report the Kullback-Leibler (KL) Divergence and the Average Marginal Wasserstein Distance between the true and modelled distributions.
@@ 3.1 Methodology
The training procedure is outlined in Algorithm 1. We first generate a static dataset of $N$ samples drawn from the prior, $\{\theta_{i}\}_{i=1}^{N}\sim\pi(\theta)$, and evaluate their corresponding likelihoods $w_{i}=L(\theta_{i})$. These likelihoods serve as importance weights. We then train a Normalizing Flow $f_{\phi}:\mathcal{Z}\to\Theta$ to map a base latent distribution $p_{Z}(z)$ to the target posterior. The network is optimized by minimizing the likelihood-weighted negative log-likelihood (see eq. 11). The workflow is shown in fig. 1.
@@ 3.2 Implementation Details
We utilize the normflows library based on PyTorch. As the loss function is architecture-agnostic, we employ the Real Non-Volume Preserving (RealNVP) architecture for all experiments. Unless otherwise stated, the base distribution $p_{Z}(z)$ is a standard multivariate Gaussian.
@@
Target Posterior KL Divergence ($D_{KL}$) Wasserstein ($W_{1,\text{avg}}$)
@@
Figure 4: Isosurface visualization of the ground truth 3-dimensional posterior distribution, displaying three distinct modes in the parameter space.
@@
Figure 5: 2D Marginal projections of the modelled 3D posterior. The flow captures the location of the modes, but exhibits connectivity artifacts due to the unimodal base distribution.
@@ 4.1 Modality of the Base Distribution
(a) $\text{Model-2D}_{1}$
(b) $\text{Model-2D}_{2}$
(c) $\text{Model-2D}_{3}$
Figure 6: Comparison of learned posteriors from Normalizing Flows with different numbers of distinct modes in the base distribution. Each sub-figure shows three 2D marginal contours corresponding to different parameter pairs.
The resulting modelled posteriors have been plotted in figs. 6 and 7 for the 2-D and 3-D toy problems, respectively. Comparing the plots, figs. 6(a) and 7(a) clearly show the artifact of connected modes. In contrast, figs. 6(b) and 7(b) exhibit partial separation, while figs. 6(c) and 7(c) achieve almost complete disconnection. This shows that if the number of modes in the true posterior is higher than that of the base distribution, then there are always tails(somewhat fat) in the modelled posterior, maintaining the topological continuity of the base distribution. One thing to keep in mind is that, although the models are different, the target posterior distribution is the same, and hence, the for all these models, the final values of the losses remain almost identical.
(a) $Model-3D_{1}$
(b) $Model-3D_{2}$
(c) $Model-3D_{3}$
Figure 7: Comparison of learned posterior from Normalizing Flows with different no. of distinct modes in the base distribution.
@@
Model W-Marginal W-Sliced KL (KDE)
@@ 4.2 Non-Gaussian Distribution
The posterior distributions considered so far are different Gaussians, either a single Gaussian or a mixture of Gaussians. In this subsection, we will look into a distribution in 3-dimensions which is a product of different non-Gaussian distributions. Given that the true posteriors are now non-Gaussian in nature and the base distributions are a mixture of Gaussians, this poses an interesting situation. We can see in fig. 8 how the modelled posterior compares with the true posterior for three different models, where, Model_iBase corresponds to the base distribution having $i$ distinct modes. The distance and divergence metrics for the three models are shown in table 3 and we can again confirm that when the modes of the base distribution and true posterior are identical, we get the best metrics.
(a) Model_1Base
(b) Model_2Base
(c) Model_3Base
Figure 8: Comparison of learned posterior from Normalizing Flows with different no. of distinct modes in the base distribution, when the true posterior is a non-Gaussian distribution.
Model W-Marginal W-Sliced KL (KDE)
Model_1Base 1.6173 1.5857 0.2426
Model_2Base 4.4237 4.1431 0.1116
Model_3Base 0.3732 0.3880 0.0940
@@ 5 Conclusion
Furthermore, we investigated the impact of the topology of the base distribution by replacing the standard Gaussian with a Gaussian Mixture Model (GMM). We observed that topological consistency is crucial: when the number of modes in the base distribution is lower than that of the target, the flow is forced to maintain connectivity, resulting in spurious “bridges” or tails between modes. Conversely, when the number of modes in the base and target distributions align, we achieve the lowest distance measures and the highest fidelity reconstruction.
@@ Acknowledgements.
The author would like to acknowledge support from ANRF (erstwhile DST-SERB), India (grant order no. CRG/2022/003208).

### v2

@@ 1 Introduction
Across diverse domains—from complex systems and finance to high-energy physics and astrophysics—scientific inquiry often relies on deriving theoretical parameters from observational data Tarantola (2005). At the core of this challenge lies the inverse problem: inferring the posterior distribution of theoretical parameters given a set of observables Cranmer et al. (2020).
@@
Traditional approaches for posterior estimation rely on sampling algorithms such as Markov Chain Monte Carlo (MCMC) Metropolis et al. (1953); Hastings (1970) and Nested Sampling (NS) Skilling (2006). In astrophysics and cosmology, implementations like emcee Foreman-Mackey et al. (2013) and dynesty Speagle (2020) have become standard tools. While these frameworks are statistically robust, they suffer significantly from the curse of dimensionality. In real-world scenarios, where the parameter space is high-dimensional and the likelihood function relies on computationally expensive simulators (e.g., in particle physics phenomenology Brehmer et al. (2018)), convergence can take weeks or even months.
@@
Recent advances in machine learning have introduced Normalizing Flows (NFs) as a powerful alternative for probabilistic modelling Tabak and Vanden-Eijnden (2010); Rezende and Mohamed (2015). By learning a bijective mapping between a simple base distribution (e.g., a Gaussian) and the complex target distribution, NFs allow for exact density estimation and efficient sampling Papamakarios et al. (2021) from the target distribution. Modern architectures, such as RealNVP Dinh et al. (2016) and Neural Spline Flows Durkan et al. (2019), offer enough expressivity to model highly complex distributions.
@@
However, standard training methodologies for NFs (typically Maximum Likelihood Estimation) rely on a critical assumption: that we possess a dataset of samples drawn from the target posterior Papamakarios et al. (2017). In the context of Simulation-Based Inference (SBI) Cranmer et al. (2020); Lueckmann et al. (2021), this often leads to approaches like Neural Posterior Estimation (NPE) Papamakarios and Murray (2016); Greenberg et al. (2019), which still require substantial simulation budgets to generate training data.
@@
In many scientific inverse problems, this assumption does not hold. We generally possess only a generic understanding of the parameters, hence assume they belong to a uniform distribution and a “black-box” simulator that can evaluate the likelihood of a given point. If we simply train a Normalizing Flow on samples drawn from the prior, the network will trivially learn to reconstruct the prior, failing to capture the information contained in the likelihood. While Variational Inference (VI) techniques exist to approximate posteriors Blei et al. (2017), applying them efficiently with flows to capture multi-modal structures remains an active area of research Wirnsberger et al. (2020); Midgley et al. (2020). Other recent efforts for estimating posterior distributions with machine learning efforts can be can be seen in Baruah et al. (2024, 2025)
@@
In this work, we propose a method to bridge this gap. By leveraging likelihood evaluations as sample weights during training—akin to Neural Importance Sampling Müller et al. (2019)—we demonstrate that a Normalizing Flow can directly learn the transformation from a simple base density to the complex target posterior, without ever seeing a “true” posterior sample. This results in an amortized inference framework that is computationally efficient and, crucially, capable of capturing the correct topological structure (modes) of the posterior distribution.
@@ 2.1 Normalizing Flows
Let $f_{\phi}:\mathcal{Z}\to\Theta$ be a differentiable bijection with a differentiable inverse (i.e., a diffeomorphism), parameterized by neural network weights $\phi$. We consider a latent space $\mathcal{Z}\subseteq\mathbb{R}^{D}$ endowed with a known, tractable base density $p_{\text{\small Z}}(z)$ (e.g., a standard Normal distribution or a Mixture of Gaussians), and a target parameter space $\Theta\subseteq\mathbb{R}^{D}$ associated with the unknown target distribution $p(\theta)$ (or the posterior $p(\theta|x)$Tabak and Vanden-Eijnden (2010); Rezende and Mohamed (2015); Papamakarios et al. (2021).
The normalizing flow $f_{\phi}$ transforms the probability mass from the base distribution to the target. Crucially, because $f_{\phi}$ is a diffeomorphism, it preserves the topological properties of the domain. Consequently, the support and connectivity of the modelled distribution $q_{\phi}(\theta)$ are strictly constrained by the topology of the base density $p_{\text{\small Z}}(z)$.
@@
$q_{\phi}(\theta)=p_{\text{\small Z}}(z)\left(z\right)\left|\det\!\left(\frac{\partial z}{\partial\theta}\right)\right|$ (1)
@@
$q_{\phi}(\theta)$ $=p_{\text{\small Z}}(z)\big(f_{\phi}^{-1}(\theta)\big)\,\left|\det\!\left(\frac{\partial f_{\phi}^{-1}(\theta)}{\partial\theta}\right)\right|$ (2)
@@
$\log q_{\phi}(\theta)$ $=\log p_{\text{\small Z}}(z)\big(f_{\phi}^{-1}(\theta)\big)+\log\left|\det\!\left(\frac{\partial f_{\phi}^{-1}(\theta)}{\partial\theta}\right)\right|$ (3)
@@
2.2 KL Divergence, Maximum Likelihood Estimation and the Loss Function
The KL divergence from a reference probability distribution $r(\theta)$ to a model distribution $q_{\phi}(\theta)$ is defined as Kullback and Leibler (1951):
$D_{\text{\small KL}}\left(r(\theta)||q_{\phi}(\theta)\right)$ $=\int r(\theta)\log\frac{r(\theta)}{q_{\phi}(\theta)}d\theta$
@@ 2.2 KL Divergence, Maximum Likelihood Estimation and the Loss Function
$\mathcal{L}(\phi)=D_{\text{\small KL}}\left(r(\theta)\|q_{\phi}(\theta)\right)-\mathbb{E}_{\theta\sim r(\theta)}\left[\log(r(\theta))\right].$ (6)
Since the entropy term $\mathbb{E}\left[\log(r(\theta))\right]$ is independent of $\phi$, minimizing the negative log-likelihood is mathematically equivalent to minimizing the forward KL divergence  Papamakarios et al. (2021):
$\min_{\phi}\left(\mathcal{L}(\phi)\right)\iff\min_{\phi}\left(D_{\text{\small KL}}\left(r(\theta)\|q_{\phi}(\theta)\right)\right).$ (7)
@@ 2.3 Application to Posterior Estimation
$D_{\text{\small KL}}\left(p(\theta\mid D)\|q_{\phi}(\theta)\right)$ $=\int p(\theta\mid D)\log\frac{p(\theta\mid D)}{q_{\phi}(\theta)}d\theta$
@@
$\min_{\phi}D_{\text{\small KL}}\left(p(\theta\mid D)\|q_{\phi}(\theta)\right)\iff\min_{\phi}\left(-\mathbb{E}_{\theta\sim p(\theta)}\left[\frac{p(D\mid\theta)}{Z}\log\left(q_{\phi}(\theta)\right)\right]\right).$ (10)
@@ 3 Implementation and Results
To evaluate the efficacy of the proposed likelihood-weighted training, we apply the method to a set of synthetic toy problems in 2 and 3 dimensions. These problems are designed with varying degrees of topological complexity (in terms of the number of modes) to test the expressivity of the flow. To quantify the fidelity of the posterior reconstruction, we report the Kullback-Leibler ($D_{\text{\small KL}}$) Divergence and the Average Marginal Wasserstein Distance between the true and modelled distributions.
@@ 3.1 Methodology
The training procedure is outlined in Algorithm 1. We first generate a static dataset of $N$ samples drawn from the prior, $\{\theta_{i}\}_{i=1}^{N}\sim\pi(\theta)$, and evaluate their corresponding likelihoods $w_{i}=L(\theta_{i})$. These likelihoods serve as importance weights. We then train a Normalizing Flow $f_{\phi}:\mathcal{Z}\to\Theta$ to map a base latent distribution $p_{\text{\small Z}}(z)$ to the target posterior. The network is optimized by minimizing the likelihood-weighted negative log-likelihood (see eq. 11). The workflow and the algorithm are shown in fig. 1.
@@ 3.2 Implementation Details
We utilize the normflows library Stimper et al. (2023) based on PyTorch. As the loss function is architecture-agnostic, we employ the Real Non-Volume Preserving (RealNVP) architecture for all experiments. Unless otherwise stated, the base distribution $p_{Z}(z)$ is a standard multivariate Gaussian.
@@
Target Posterior KL Divergence ($D_{\text{\small KL}}$) Wasserstein ($W_{1,\text{avg}}$)
@@
Figure 4: Isosurface visualization of the ground truth 3-dimensional posterior distribution, displaying three distinct modes in the parameter space. The vertical axes share identical limits and tick spacing across all panels.
@@
Figure 5: 2D Marginal projections of the modelled 3D posterior. The flow captures the location of the modes, but exhibits connectivity artifacts due to the unimodal base distribution. The vertical axes share identical limits and tick spacing across all panels
@@ 4.1 Modality of the Base Distribution
Figure 6: Comparison of learned posteriors from Normalizing Flows with different numbers of distinct modes in the base distribution. From Left to Right, the models have 1 base mode($Model-2D_{1}$), 2 base modes($Model-2D_{2}$) and 3 base modes($Model-2D_{3}$).
The resulting modelled posteriors have been plotted in figs. 6 and 7 for the 2-D and 3-D toy problems, respectively. Comparing the plots, the modelled samples from $Model-2D_{1}$ and $Model-3D_{1}$ figs. 6 and 7 clearly show the artifact of connected modes. In contrast, samples from $Model-2D_{2}$ and $Model-3D_{2}$ figs. 6 and 7 exhibit partial separation, while $Model-2D_{3}$ and $Model-3D_{3}$ figs. 6 and 7 achieve almost complete disconnection. This shows that if the number of modes in the true posterior is higher than that of the base distribution, then there are always tails(somewhat fat) in the modelled posterior, maintaining the topological continuity of the base distribution. One thing to keep in mind is that, although the models are different, the target posterior distribution is the same, and hence, the for all these models, the final values of the losses remain almost identical.
Figure 7: Comparison of learned posterior from Normalizing Flows with different no. of distinct modes in the base distribution. From Left to Right, the models have 1 base mode($Model-3D_{1}$), 2 base modes($Model-3D_{2}$) and 3 base modes($Model-3D_{3}$). The top and bottom rows are the projections on x-y plane and x-z plane, respectively.
@@
Model Wasserstein ($W_{1,\text{avg}}$) W-Sliced KL Divergence ($D_{\text{\small KL}}$)
@@ 4.2 Non-Gaussian Distribution
The posterior distributions considered so far are different Gaussians, either a single Gaussian or a mixture of Gaussians. In this subsection, we will look into a distribution in 3-dimensions which is a product of different non-Gaussian distributions. Given that the true posteriors are now non-Gaussian in nature and the base distributions are a mixture of Gaussians, this poses an interesting situation. We can see in fig. 8 how the modelled posterior compares with the true posterior for three different models, where, Model-nonGaussi corresponds to the base distribution having $i$ distinct modes. The distance and divergence metrics for the three models are shown in table 3 and we can again confirm that when the modes of the base distribution and true posterior are identical, we get the best metrics.
Figure 8: Comparison of learned posterior from Normalizing Flows with different no. of distinct modes in the base distribution, when the true posterior is a non-Gaussian distribution. From Left to Right, the models have 1 base mode(Model-nonGauss1), 2 base modes(Model-nonGauss2) and 3 base modes(Model-nonGauss3). The top and bottom rows are the projections on x-y plane and x-z plane, respectively.
Model Wasserstein ($W_{1,\text{avg}}$) W-Sliced KL Divergence ($D_{\text{\small KL}}$)
Model-nonGauss1 1.6173 1.5857 0.2426
Model-nonGauss2 4.4237 4.1431 0.1116
Model-nonGauss3 0.3732 0.3880 0.0940
@@ 5 Conclusion
Furthermore, we investigated the impact of the topology of the base distribution by replacing the standard Gaussian with a Gaussian Mixture Model (GMM). We observed that topological consistency is crucial: when the number of modes in the base distribution is lower than that of the target, the flow is forced to maintain connectivity, resulting in spurious “bridges” or tails between modes. Conversely, when the number of modes in the base and target distributions align, we achieve the lowest distance measures and the highest fidelity reconstruction. Further discussions about topology-aware generative algorithms can be found in  Ngairangbam et al. (2025); Jang et al. (2019).
@@ Acknowledgements.
The author would like to acknowledge support from ANRF (erstwhile DST-SERB), India (grant order no. CRG/2022/003208). The author thanks Sunando Kumar Patra, Bangabasi Evening College, and Subhadeep Mondal, Bennett University, for their helpful insights and guidance on the analysis of distribution topology and reconstruction fidelity, and in preparing this manuscript.